# Towards Understanding the Robustness of Diffusion-Based Purification: A Stochastic Perspective

### v2

@@ 4.2 Explaining Stochasticity-Driven Robustness
Table 1: Evaluation of state-of-the-art DBP methods, EoT significantly influences the evaluation accuracy (%) of model robustness.
DiffPure GDMP (MSE) DPDDPM DPDDIM
@@
PGD20-EoT1 69.04 53.13 60.25 54.59
PGD20-EoT10 55.96 40.97 47.27 42.19
To elucidate the robustness of DBP models, particularly under EoT evaluations, we analyze the performance of several DBP models—DiffPure, GDMP, DPDDPM, and DPDDIM—under white-box attacks with and without Expectation over Transformation (EoT) iterations (denoted as EoT10 and EoT1, respectively). The results shown in Table 1 suggest that these DBP models remain robust under white-box attacks, with EoT evaluations resulting in moderate reductions in robustness. The detailed discussion on the selection of PGD and EoT steps is provided in Appendix E and Appendix A.
To gain deeper insights, we visualize the attack trajectories using t-SNE, projecting them onto an xy-plane with loss values represented along the z-axis. We compare trajectories for three types of attacks: white-box without EoT (White-box), white-box with EoT (White-box-EoT), and Deterministic White-box (DW-box). As shown in Figure 1, the trajectories exhibit high variance across all settings, reflecting the stochastic nature of DBP models. Specifically, DW-box attacks lead to a significant increase in loss values, whereas white-box attacks, even with EoT, result in only moderate increases. This suggests that stochasticity prevents attackers from finding the optimal attack direction. Specifically, due to the significant variance of the attack gradients, even if the EoT direction is an accurate estimation of the mean gradient direction, it may not be completely consistent with the most effective direction corresponding to the DW-box attack, thus resulting in a decline in attack performance. Additional evidences are provided in Appendix B.
@@
Further analysis of the loss landscape, presented in Figure 3, illustrates key differences between White-box-EoT and Deterministic White-box attacks. The trajectory of the White-box-EoT attack diverges from the Deterministic White-box direction, resulting in a flatter loss landscape along the White-box-EoT path. This behavior indicates that White-box-EoT attacks fail to identify the most effective direction due to the stochastic nature of DBP models. In contrast, the Deterministic White-box attack induces a sharp increase in loss, revealing that when stochasticity is removed, the model becomes more vulnerable to adversarial perturbations. These findings differ from models trained using AT, where the whole loss landscape tends to remain flat and resistant across adversarial directions (Shafahi et al., 2019).
To conclude, it is suggested that instead of possessing a flat loss landscape, DBP models rely on stochasticity to evade the most effective attack directions. Note that while certified defense methods like random smoothing also incorporate stochasticity (Xiao et al., 2022; Carlini et al., 2022), their mechanisms and implications differ from those of DBP methods, as discussed in Appendix D.
5 Towards Improving the Purification Capability of DBP
Based on the analysis from Section 4, although the stochasticity-driven robustness of DBP does not depend on the flatness of the loss landscape, flattening the landscape can still benefit the DBP robustness given the non-trivial loss increment along the EoT direction. To achieve a flat loss landscape, we need to introduce adversarial samples to the training of the DBP models and minimize the loss on them. From the perspective of adversarial purification, this amounts to improving the diffusion model in its ability to purify adversarial perturbations.
To this end, we propose Adversarial Denoising Diffusion Training (ADDT), which integrates adversarial perturbations into the training of the diffusion model in DBP. ADDT employs an iterative two-step procedure: (1) Classifier-Guided Perturbation Optimization (CGPO), which generates adversarial perturbations by maximizing the classification error of a pre-trained classifier; (2) Diffusion Model Training, which updates the diffusion model using these perturbations to improve its capability of adversarial purification.
Integrating adversarial perturbations into diffusion training poses a challenge due to the Gaussian noise assumption inherent in diffusion models. To address this, we introduce Rank-Based Gaussian Mapping (RBGM), a technique designed to transform adversarial perturbations into a form consistent with the Gaussian noise assumption. RBGM renders the perturbations more “Gaussian-like”, facilitating their integration into the diffusion training process.
An overview of ADDT is illustrated in Figure 4, with pseudocode in Appendix G. The following subsections detail the components of ADDT.
@@ 5.1 Adversarial Denoising Diffusion Training
Classifier-Guided Perturbation Optimization (CGPO) step. In this step, we aim to refine adversarial perturbations $\bm{\delta}$ in a way that maximizes the classification error of a pre-trained classifier $C$. The process starts by reconstructing a clean image $\hat{\bm{x}}_{0}$ from the perturbed input $\bm{x}^{\prime}_{t}$ using the diffusion model $\bm{P}$. $\bm{P}(\bm{\theta},\bm{x}^{\prime}_{t},t)$ denotes a one-step diffusion process, which takes the noisy input $\bm{x}^{\prime}_{t}$ and time step $t$ and reconstructs the image $\hat{\bm{x}}_{0}$, following the formulation in Equation 4. The classifier $C$ is then applied to this reconstructed image $\hat{\bm{x}}_{0}$ to predict a label. To maximize the prediction error compared to the true label $y$, the optimization objective for refining $\bm{\delta}$ can be defined as:
$\bm{\delta}^{*}=\arg\max_{\bm{\delta}}\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\left[L\left(C\left(\bm{P}\left(\bm{\theta},\bm{x}^{\prime}_{t}(\bm{x}_{0},\bm{\epsilon},\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})),t\right)\right),y\right)\right],$ (6)
where $L(\cdot,y)$ denotes the loss function used to measure the discrepancy between the classifier’s predicted label and the true label $y$. During the optimization, since RBGM is non-differentiable, we accumulate the gradient $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$ to $\bm{\delta}$. Notably, the classifier in this process serves purely for semantic guidance and does not have to be consistent with the protected model. Further discussions on cross-classifier performance are provided in Table 5.
Diffusion Model Training step. The goal of this step is to update the diffusion model parameters to accurately recover the original image $\bm{x}_{0}$ from a perturbed version $\bm{x}^{\prime}_{t}$. As depicted on the right side of Figure 4, The model is optimized to subtract both the Gaussian noise and the RBGM-mapped adversarial perturbations, effectively denoising the input. The optimization objective is defined as:
$\bm{\theta}^{*}=\arg\min_{\bm{\theta}}\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\left[\frac{\sqrt{\overline{\alpha}_{t}}}{\sqrt{1-\overline{\alpha}_{t}}}\left\|\bm{x}_{0}-\bm{P}(\bm{\theta},\bm{x}^{\prime}_{t},t)\right\|_{2}^{2}\right],$ (7)
where the expectation is taken over the distribution of original images $\bm{x}_{0}\sim\mathcal{D}$, time steps $t\sim\mathcal{U}(\{1,\dots,T\})$, and Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$. The perturbed input $\bm{x}^{\prime}_{t}$ is formed from the original image $\bm{x}_{0}$, Gaussian noise $\bm{\epsilon}$, and the RBGM-mapped adversarial perturbation $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$, as defined in Equation 8. The scaling factor ${\sqrt{\overline{\alpha}_{t}}}/{\sqrt{1-\overline{\alpha}_{t}}}$ ensures consistency with the standard formulation of DDPM/DDIM loss. This factor reflects the expected squared error between the noise introduced to the input and the noise removed by the diffusion model during denoising.
@@ 5.2 Rank-Based Gaussian Mapping
Traditional diffusion models operate under the premise that input images are corrupted by independent Gaussian noise $\bm{\epsilon}$. To ensure that the perturbations remain Gaussian-like while capturing adversarial characteristics, we introduce the Rank-Based Gaussian Mapping (RBGM), illustrated in Figure 5.
Figure 5: Rank-Based Gaussian Mapping. RBGM trims the input to follow Gaussian distribution. It samples a Gaussian noise and then replaces elements in the input with those from the Gaussian noise, matched according to their respective ranks.
The RBGM function, denoted by $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$, takes a perturbation $\bm{\delta}$ as input. The key idea is to preserve the rank ordering of the elements in $\bm{\delta}$ but replace their actual values with those from a standard Gaussian distribution. Specifically, we sample a Gaussian tensor $\bm{\epsilon}_{s}$ of the same dimensions as $\bm{\delta}$. We then sort the elements of both $\bm{\delta}$ and $\bm{\epsilon}_{s}$ respectively in ascending order. By mapping the sorted elements of $\bm{\delta}$ to the corresponding elements of $\bm{\epsilon}_{s}$, we obtain $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$, which approximates Gaussian-distributed but retains the structural information of $\bm{\delta}$. To further enhance the Gaussian nature of the noise, we mix the RBGM-mapped perturbation with additional random Gaussian noise.
By combining the RBGM-induced perturbation with Gaussian noise, we generate an adversarial input $\bm{x}^{\prime}_{t}$ as follows:
@@
where $\lambda_{t}$ modulates the level of adversarial perturbation. This ensures that the overall noise remains largely independent of $\bm{x}_{0}$ and that the perturbations do not overwhelm the denoising model’s learning capabilities. We determine $\lambda_{t}$ using the following formulation:
$\lambda_{t}=\texttt{clip}(\gamma_{t}\lambda_{unit},\lambda_{min},\lambda_{max}),\quad\gamma_{t}=\frac{\sqrt{\overline{\alpha}_{t}}}{\sqrt{1-\overline{\alpha}_{t}}},$ (9)
where the clip function limits $\lambda_{t}$ between $\lambda_{min}$ and $\lambda_{max}$.
@@ 6 Experiments and Discussions
6.1 Experiment Setups
@@
Diffusion model DBP model $Clean$ $l_{\infty}$ $l_{2}$
@@
Table 3: Clean and robust accuracy (%) on DPDDPM. ADDT improve robustness across different NFE, especially at lower NFE (*: default DDPM generation setting; -: classifier only).
Dataset NFE Vanilla ADDT
$Clean$ $l_{\infty}$ $l_{2}$ $Clean$ $l_{\infty}$ $l_{2}$
@@
Classifier. We train a WideResNet-28-10 for 200 epochs following the methods in (Yoon et al., 2021; Wang et al., 2022), achieving $95.12\%$ accuracy on CIFAR-10 and $76.66\%$ on CIFAR-100 dataset.
DBP timestep. For the diffusion forward process, we adopt the same timestep settings as DiffPure (Nie et al., 2022). In continuous-time models, such as the VPSDE (DDPM++) variant, with the forward time parameter $0\leq t\leq 1$, we set $t^{*}=0.1$, which strikes a balance between noise introduction and computational efficiency. For discrete-time models, such as DDPM and DDIM, where $t=0,1,...,T$, we similarly set the timestep to $t^{*}=0.1\times T$. Additional settings and results on DPEDM are provided in Appendix I.
Model Classifier Vanilla ADDT
$Clean$ $l_{\infty}$ $l_{2}$ $Clean$ $l_{\infty}$ $l_{2}$
DPDDPM-1000 VGG-16 (Simonyan & Zisserman, 2014) 84.77 41.99 66.89 85.06 46.09 67.87
ResNet-50 (He et al., 2016) 83.11 44.04 67.58 83.84 48.14 67.87
ViT-B (Dosovitskiy et al., 2020) 85.45 45.61 69.53 85.25 48.63 69.92
DPDDIM-100 VGG-16 (Simonyan & Zisserman, 2014) 87.16 29.00 61.82 87.55 35.06 66.11
ResNet-50 (He et al., 2016) 86.04 31.74 62.11 86.57 38.77 65.82
ViT-B (Dosovitskiy et al., 2020) 88.77 34.38 65.72 88.48 41.02 68.65
Robustness evaluation. We employ PGD20+EoT10 (Athalye et al., 2018) for assessing model robustness. For $\ell_{\infty}$-norm attacks, we set the step size $\alpha=2/255$ and the maximum perturbation $\epsilon=8/255$, while for $\ell_{2}$-norm attacks, we use $\alpha=0.1$ and $\epsilon=0.5$. Due to the high computational cost of EoT attacks, we evaluate our models on the first 1024 images for CIFAR-10 and CIFAR-100 datasets.
ADDT. ADDT fine-tuning is guided by the pre-trained WideResNet-28-10 classifier. For the CIFAR-10 dataset, we utilize the pre-trained exponential moving average (EMA) diffusion model from Ho et al. (2020), which has been converted into the Huggingface Diffusers format by Fang et al. (2023). For the CIFAR-100 dataset, we fine-tune this CIFAR-10 diffusion model over 100 epochs. In CGPO, we set the hyperparameters to $\lambda_{unit}=0.03$, $\lambda_{min}=0$, and $\lambda_{max}=0.3$, and iteratively refine the perturbation $\delta$ for 5 steps. Additional details regarding computational cost are provided in Appendix O.

### v3

@@ 4.2 Explaining Stochasticity-Driven Robustness
When attacking stochastic models, a commonly used technique is EoT (Athalye et al., 2018b). To assess the influence of adaptive attacks, we compare the performance of white-box attacks with and without using EoT. The results in Table 1 show that EoT-based adaptive attacks have only a modest impact on robustness, which contrasts sharply with DW-box attacks. A more detailed discussion of EoT steps is provided in Appendix A.
Table 1: Evaluation of DBP methods under various attack settings shows that EoT significantly affects the model’s robustness accuracy (%).
Metric DiffPure GDMP (MSE) DPDDPM DPDDIM
@@
PGD20 69.04 53.13 60.25 54.59
PGD20+EoT10 55.96 40.97 47.27 42.19
To compare different attacks, we visualize the attack trajectories using t-SNE. These trajectories are projected onto the xy-plane, with loss values plotted on the z-axis. We compare three types of attacks: white-box without EoT (White-box), white-box with EoT (White-box-EoT), and Deterministic White-box (DW-box). As shown in Figure 1, the trajectories vary significantly across all settings, reflecting the stochasticity of DBP models. Notably, DW-box attacks lead to a significant increase in loss values, whereas white-box attacks—even with EoT—result in only moderate increases. This observation suggests that stochasticity prevents attackers from finding the optimal attack direction. Even if the EoT estimate accurately captures the mean gradient direction, the high variance in attack gradients may prevent alignment with the optimal attack direction (the direction of the DW-box attack). This leads to reduced attack performance. Additional evidence is provided in Appendix B.
@@
To study how different attacks impact robust accuracy, we compared the loss landscapes under White-box-EoT and Deterministic White-box attacks. We computed the average loss variation across a batch of images and present the results in Figure 3. The trajectory of the White-box-EoT attack deviates from that of the Deterministic White-box attack, resulting in a flatter loss landscape along its path, whereas the Deterministic White-box attack produces a steep increase in loss. This observation suggests that the inherent stochasticity in DBP models prevents White-box-EoT attacks from identifying the optimal attack direction, and that removing this stochasticity makes the model vulnerable to adversarial perturbations. These findings contrast with adversarially trained models, where the loss landscape remains flat even along adversarial directions (Shafahi et al., 2019). Further details on the loss landscape visualization can be found in Appendix F.
In conclusion, we propose that DBP models, instead of exhibiting a flat loss landscape, leverage stochasticity to evade the most effective attack directions. Note that while certified defense methods such as random smoothing also incorporate stochasticity (Xiao et al., 2022; Carlini et al., 2022), their mechanisms and implications differ from those of DBP methods (see Appendix D).
5 Towards Improving the Purification Capability of DBP Models
Based on the analysis in Section 4, and considering the loss increase along the DW-box attack direction, we propose that as an alternative to stochasticity-driven robustness, the performance of DBP can be further improved by flattening the loss landscape. Achieving this requires incorporating adversarial samples into the training of DBP models. From the perspective of adversarial purification, this involves enhancing the diffusion model’s ability to purify adversarial perturbations.
To address this, we propose Adversarial Denoising Diffusion Training (ADDT), a method that incorporates adversarial perturbations into the training of diffusion models in DBP. ADDT follows an iterative two-step process: (1) Classifier-Guided Perturbation Optimization (CGPO), which generates adversarial perturbations by maximizing the classification error of a pre-trained classifier; (2) Diffusion Model Training, which trains the diffusion model on these perturbations to improve its ability to purify adversarial perturbations.
Integrating adversarial perturbations into diffusion training is challenging due to the Gaussian noise assumption inherent in diffusion models. To overcome this, we propose Rank-Based Gaussian Mapping (RBGM), a technique that transforms adversarial perturbations into a form more consistent with the Gaussian noise assumption, thereby facilitating their integration into the diffusion training process.
Figure 4 provides an overview of ADDT, and pseudocode is available in Appendix G. The following subsections detail the components of ADDT.
@@ 5.1 Adversarial Denoising Diffusion Training
Classifier-Guided Perturbation Optimization (CGPO) step. In this step, we refine the adversarial perturbations $\bm{\delta}$ to maximize the classification error of a pre-trained classifier $C$, as shown on the left side of Figure 4. The process begins by reconstructing a clean image $\hat{\bm{x}}_{0}$ from the perturbed input $\bm{x}^{\prime}_{t}$ using the diffusion model $\bm{P}$. The perturbed input $\bm{x}^{\prime}_{t}$ is generated from the original image $\bm{x}_{0}$, Gaussian noise $\bm{\epsilon}$, and the RBGM-mapped perturbation $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$, as detailed in Section 5.2. Here, $\bm{P}(\bm{\theta},\bm{x}^{\prime}_{t},t)$ represents a one-step diffusion process based on the formulation in Equation 4. It reconstructs the image $\hat{\bm{x}}_{0}$ from the noisy input $\bm{x}^{\prime}_{t}$. The classifier $C$ is then used to predict a label for this reconstructed image $\hat{\bm{x}}_{0}$. The optimization objective is to maximize the discrepancy between the classifier’s prediction and the ground-truth label $y$. This can be formulated as:
$\bm{\delta}^{*}=\operatorname{arg\,max}_{\bm{\delta}}\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\left[L\left(C\left(\bm{P}\left(\bm{\theta},\bm{x}^{\prime}_{t}(\bm{x}_{0},\bm{\epsilon},\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})),t\right)\right),y\right)\right],$ (6)
where $L(\cdot,y)$ denotes the loss function measuring the discrepancy between the classifier’s prediction and the ground-truth label $y$. The pre-trained classifier $C$ provides semantic guidance and need not be identical to the protected classifier, as detailed in Section 6.2. To optimize $\bm{\delta}$, we employ the PGD attack (Madry et al., 2018). Since RBGM is non-differentiable, we apply the Backward Pass Differentiable Approximation (BPDA) approach (Athalye et al., 2018a). By applying BPDA, we employ RBGM during forward propagation but bypass it during backpropagation. Note that BPDA/RBGM is only applied during training and does not affect the evaluation of DBP robustness.
Diffusion Model Training step. The objective of this step is to update the diffusion model’s parameters so that it can reconstruct the original image $\bm{x}_{0}$ from its perturbed counterpart $\bm{x}^{\prime}_{t}$. As illustrated on the right side of Figure 4, the model aims to remove both Gaussian noise and adversarial perturbations, effectively denoising the input. The optimization objective is defined as:
$\bm{\theta}^{*}=\operatorname{arg\,min}_{\bm{\theta}}\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\left[\frac{\sqrt{\overline{\alpha}_{t}}}{\sqrt{1-\overline{\alpha}_{t}}}\left\|\bm{x}_{0}-\bm{P}(\bm{\theta},\bm{x}^{\prime}_{t}(\bm{x}_{0},\bm{\epsilon},\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})),t)\right\|_{2}^{2}\right],$ (7)
where the scaling factor ${\sqrt{\overline{\alpha}_{t}}}/{\sqrt{1-\overline{\alpha}_{t}}}$ ensures consistency with the standard DDPM/DDIM loss formulation.
@@ 5.2 Rank-Based Gaussian Mapping
Traditional diffusion models assume that input images are corrupted by independent Gaussian noise $\bm{\epsilon}$. To preserve Gaussian-like characteristics while encoding adversarial features, we introduce Rank-Based Gaussian Mapping (RBGM), as shown in Figure 5. RBGM, denoted as $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$, retains the ordering of the elements in the input tensor $\bm{\delta}$ while replacing their original values with samples drawn from a standard Gaussian distribution $\bm{\epsilon}_{s}$. Specifically, we first sample a Gaussian tensor $\bm{\epsilon}_{s}$ with dimensions matching $\bm{\delta}$. Next, we sort the elements of both $\bm{\delta}$ and $\bm{\epsilon}_{s}$ in ascending order. By mapping the sorted elements of $\bm{\delta}$ to the corresponding elements of $\bm{\epsilon}_{s}$, we obtain $\bm{\epsilon}_{\bm{\delta}}(\bm{\delta})$, which approximates a Gaussian distribution while preserving the structural information in $\bm{\delta}$. To further refine the perturbation towards Gaussianity, we blend the RBGM-mapped perturbation with additional randomly sampled Gaussian noise. The adversarial input $\bm{x}^{\prime}_{t}$ is then defined as follows:
@@
where $\lambda_{t}$ modulates the level of adversarial perturbation. This ensures that the overall perturbation remains largely independent of $\bm{x}_{0}$, preventing the perturbations from overwhelming the denoising model’s learning process. We determine $\lambda_{t}$ using the following formulation:
$\lambda_{t}=\texttt{clip}(\gamma_{t}\lambda_{unit},\lambda_{min},\lambda_{max}),\quad\gamma_{t}=\sqrt{\overline{\alpha}_{t}}/\sqrt{1-\overline{\alpha}_{t}},$ (9)
where the clip function constrains $\lambda_{t}$ between $\lambda_{min}$ and $\lambda_{max}$. Additional details and discussions about RBGM can be found in Appendix K.
@@ 6 Experiments and Discussions
6.1 Experiment Setup
Figure 5: Rank-Based Gaussian Mapping. RBGM trims the input to follow Gaussian distribution. It samples a Gaussian noise and then replaces elements in the input with corresponding values from the noise, based on matching ranks.
Classifier. We train a WideResNet-28-10 model for 200 epochs following the methods in Yoon et al. (2021); Wang et al. (2022), achieving an accuracy of $95.12\%$ on CIFAR-10 and $76.66\%$ on CIFAR-100 dataset.
DBP timestep. For the diffusion forward process, we adopt the same timestep settings as DiffPure (Nie et al., 2022). In continuous-time models, such as the VPSDE (DDPM++) variant, with the forward time parameter $0\leq t\leq 1$, we set $t^{*}=0.1$, balancing noise introduction and computational efficiency. For discrete-time models, such as DDPM and DDIM, where $t=0,1,...,T$, we similarly set the timestep to $t^{*}=0.1\times T$. Additional settings and results for DPEDM are provided in Appendix I.
DBP method Classifier Vanilla ADDT
$Clean$ $\ell_{\infty}$ $\ell_{2}$ $Clean$ $\ell_{\infty}$ $\ell_{2}$
DPDDPM VGG-16 (Simonyan & Zisserman, 2014) 84.77 41.99 66.89 85.06 46.09 67.87
ResNet-50 (He et al., 2016) 83.11 44.04 67.58 83.84 48.14 67.87
ViT-B (Dosovitskiy et al., 2020) 85.45 45.61 69.53 85.25 48.63 69.92
DPDDIM VGG-16 (Simonyan & Zisserman, 2014) 87.16 29.00 61.82 87.55 35.06 66.11
ResNet-50 (He et al., 2016) 86.04 31.74 62.11 86.57 38.77 65.82
ViT-B (Dosovitskiy et al., 2020) 88.77 34.38 65.72 88.48 41.02 68.65
Robustness evaluation. We assess model robustness using the PGD20+EoT10 attack (Athalye et al., 2018b). For $\ell_{\infty}$-norm attacks, we set the step size $\alpha=2/255$ and the maximum perturbation $\epsilon=8/255$; for $\ell_{2}$-norm attacks, we use $\alpha=0.1$ and $\epsilon=0.5$. Due to the high computational cost of EoT attacks, we evaluate the models on the first 1024 images from the CIFAR-10 and CIFAR-100 datasets. For comprehensive evaluation settings, including an analysis of AutoAttack (Croce & Hein, 2020) performance on DBP models, refer to Appendix E.
ADDT. ADDT fine-tuning is guided by the pre-trained WideResNet-28-10 classifier. For the CIFAR-10 dataset, we utilize the pre-trained exponential moving average (EMA) diffusion model developed by Ho et al. (2020) (converted to Huggingface Diffusers format by Fang et al. (2023)). For the CIFAR-100 dataset, we fine-tune the CIFAR-10 diffusion model for 100 epochs. In CGPO, we set the hyperparameters to $\lambda_{unit}=0.03$, $\lambda_{min}=0$, and $\lambda_{max}=0.3$, and iteratively refine the perturbation $\bm{\delta}$ for 5 steps. Additional details regarding computational cost are provided in Appendix O.
@@ 6.1 Experiment Setup
Diffusion model DBP method $Clean$ $\ell_{\infty}$ $\ell_{2}$
@@
Table 4: Clean and robust accuracy (%) on DPDDPM. ADDT improves robustness across different NFEs, especially at lower NFEs (*: default DDPM generation setting; -: classifier only).
Dataset NFEs Vanilla ADDT
$Clean$ $\ell_{\infty}$ $\ell_{2}$ $Clean$ $\ell_{\infty}$ $\ell_{2}$